OpenStudy (anonymous):

What is the 8th term of the geometric sequence where a1 = 243 and a4 = 9?

OpenStudy (anonymous):

243 ,243r,243r^2,9 243*r^3=9 r=1/3

OpenStudy (anonymous):

As the magnitude of the numbers decreases, its safe to assume you are dealing with division or subtraction. Division would say: 243 /x =a2 a2/x = a3 a3/x = 9

OpenStudy (anonymous):

so can u find now? @lkiamx

OpenStudy (anonymous):

where did the 243 go?

OpenStudy (anonymous):

also, how did u get 1/3??

OpenStudy (anonymous):

243 is the first term 243/3=81 81/3=27 27/3=9 bcz r=1/3 ok?

OpenStudy (anonymous):

@lkiamx is it clear?

OpenStudy (anonymous):

Every new step of the sequence, you divide by 3 as sorour answered.

OpenStudy (anonymous):

@sorour yes :)

OpenStudy (anonymous):

@TimSmit okay, I get it now

OpenStudy (anonymous):

@lkiamx good

OpenStudy (anonymous):

Thanks for helping me :) have a nice day

OpenStudy (anonymous):

ur welome,nice dayfor u too :)

OpenStudy (anonymous):

You too :)